SLAC PEP-II
BABAR
SLAC<->RAL
Babar logo
HEPIC E,S & H Databases PDG HEP preprints
Organization Detector Computing Physics Documentation
Personnel Glossary Sitemap Search Hypernews
Unwrap page!
Det. Search
Who's who?
Meetings
FAQ
Images
Archive
Systems
Performance
Intern. region
Vertex Tracker
Drift chamber
DIRC
Calorimeter
IFR
LST
Magnet
Electronics
Trigger
Operations
Run Coordination
Contact Experts
Shift Takers Info
Operations Manual
Electronic Logbook
Ops Hypernews
Shift Signup
Check this page for HTML 4.01 Transitional compliance with the
W3C Validator
(More checks...)

Minutes of the Trigger Operation Meeting (Dec.6, 2005)

Attendees: Al Eisner, Fang Fang, Selina Li, Rahmat Rahmat, Dong Su, Dan Walker.

Operations

  • The Trigger parallel session for the Dec. Collaboration meeting will be next Tuesday morning. Each trigger subsystem expert should be prepared to give a report.
  • Rahmat is officially the QA manager for Trigger.
  • We are continuing to take good data. Luminosity is getting higher and background condition also improves. There are some down time due to PEP.

EMT

  • Dan showed some plots for the EMT turn-on curves. He will re-organize these plots and select the important ones to use for OPR.
  • We will change the prescale factor to be greater than 1 as soon as the next online release is installed at IR2.
  • Selina noticed a dip in theta=0, phi=12 region for runs in Run 5b. This corresponds to masked EMC channels. Also a hot tower was seen in theta=4, phi=28 region in runs 59751-59760 taken on Nov.30. Dan found out that this hot tower was caused by a noisy ADB. This was masked off by the EMC when the problem re-occurred on Fri 02 Dec, around run 59842. Since then this tower is empty on the EMT plots. Rahmat will also check more carefully with these runs after they are processed.

DCT/DCZ

  • The TSF firmware testing continues at the test stand. Last week, Fang, Su Dong and Selina got the test stand set up to run Xuedong's interface test and used an oscilloscope to check if the frame bit signals were there or not. The frame bits were there even after the fast control cable was pulled and then plugged back in. As Marc pointed out that the DCM clock should be reset; however, we found that only the Y board did and the X boards weren't reset. This might explain the vulnerability seen in the X boards but not the Y boards at IR2. We will need to double check the header to see if reset was done. Double peaks were also seen in the Y board. This needs to be understood better at the test stand.

GLT

  • Some physics MC samples were generated, and Su Dong will have some line configuration information by the Dec. collaboration meeting.

Online

  • Xuedong's interface code was disappeared at some point, and Fang got it recovered to use for the interface test with the new TSF firmware.
  • Fang has been studying the L1DctDiagOnline and L1DctOdf packages to separate codes for test stand and online use. Maybe we would have different versions of the library for calibration and online.

L3

  • The L3 reference histogram disappeared and got fixed. See HN posted, its link and follow-ups. Al suggested that we might want to update the reference file for L3 after the running condition becomes more stable.
  • We got paged for L3 damage over this weekend. It turned out to be a radiation upset problem with the DCH electronics. The Trigger long-term instructions for the pilot and DQM are updated to show the shifters how to help quickly diagnose if it was a trigger problem or other subsystem's.

Data Quality and Monitoring

  • Rahmat is performing the normal QA checks for Run4 reprocessing. Nothing unusual was seen.
  • Rahmat showed some qualities in the strip charts for OPR QA checks on the runs from run5b, and asked what else should be added to the strip charts. Any requests and suggestions should be sent to Rahmat.

This page is created by Selina Li (selina@slac.stanford.edu); however, you are welcome to edit and/or correct above items, specially for those who attended the meeting. Thanks.